# Sharper Exponential Convergence Rates for Sinkhorn’s Algorithm in Continuous Settings

### v1

@@ Abstract
We study the convergence rate of Sinkhorn’s algorithm for solving entropy-regularized optimal transport problems when at least one of the probability measures, $\mu$, admits a density over $\mathbb{R}^{d}$. For a semi-concave cost function bounded by $c_{\infty}$ and a regularization parameter $\lambda>0$, we obtain exponential convergence guarantees on the dual sub-optimality gap with contraction rate polynomial in $\lambda/c_{\infty}$. This represents an exponential improvement over the known contraction rate $1-\Theta(\exp(-c_{\infty}/\lambda))$ achievable via Hilbert’s projective metric. Specifically, we prove a contraction rate value of $1-\Theta(\lambda^{2}/c_{\infty}^{2})$ when $\mu$ has a bounded log-density. In some cases, such as when $\mu$ is log-concave and the cost function is $c(x,y)=-\langle x,y\rangle$, this rate improves to $1-\Theta(\lambda/c_{\infty})$. The latter rate matches the one that we derive for the transport between isotropic Gaussian measures, indicating tightness in the dependency in $\lambda/c_{\infty}$. Our results are fully non-asymptotic and explicit in all the parameters of the problem.
@@ 1 Introduction
$\sup_{\phi\in L^{1}(\mu),\psi\in L^{1}(\nu)}\int\phi\mathrm{d}\mu+\int\psi\mathrm{d}\nu+\lambda\left(1-\int\int\exp\left(\frac{\phi\oplus\psi-c}{\lambda}\right)\mathrm{d}\mu\mathrm{d}\nu\right),$ (3)
where $(\phi\oplus\psi)(x,y)=\phi(x)+\psi(y)$. The classical approach to solve (3) is via Sinkhorn’s algorithm. Denoting $F(\phi,\psi)$ the dual functional being maximized in (3), Sinkhorn’s algorithm starts at an arbitrary point $\psi_{0}\in L^{1}(\nu)$ and repeatedly performs exact maximization of the objectives $\phi\mapsto F(\phi,\psi)$ and $\psi\mapsto F(\phi,\psi)$. This leads to the sequence of iterates
@@ 1.1 Comparison with related work
Iterations (4), which we refer to as Sinkhorn’s algorithm, were rediscovered several times and perhaps first appeared in [30]. Consequently, these iterations and close variants thereof are known under various names in various communities, such as the Sinkhorn-Knopp algorithm, Iterative Proportional Fitting Procedure (IPFP), RAS algorithm, matrix scaling algorithm, Bregman algernative projection or Fortet’s iterations; see the surveys [19, 25] for some historical remarks. Most quantitative convergence analyses can be classified into two categories: (i) exponential convergence rates with non-robust constants and (ii) polynomial convergence rates with robust constants.
Convergence guarantees of the first kind can be obtained by noticing that Sinkhorn’s iteration is a contraction for Hilbert’s projective metric [16] where the contraction rate can be shown to be at most $1-\Theta(e^{-c_{\infty}/\lambda})$; see [9] for a recent extension to non-compact settings and [25, Remark 4.12] for a bibliographical discussion. As put forth in [25, Remark 4.15], the dependence on $c_{\infty}/\lambda$ of this contraction rate numerically seems tight in the most difficult cases, e.g. when the cost $c$ is random. In more favorable cases however, such as when there exists an optimal transport map solving problem (1) with $\lambda=0$, many have observed experimentally that the dependency in $c_{\infty}/\lambda$ is generally much better in practice.
Convergence guarantees of the second kind can be traced back to [21], and have more recently regained interest following [1]. In their tightest form, due to [15], one has the guarantee $\delta_{t}\leq 2c_{\infty}^{2}/(\lambda t)$ for $t\geq 1$ (see [8, Proposition 10]). This result relies on an explicit expression (see Eq. (25)) for the one-step dual improvement $\delta_{t}-\delta_{t+1}$, which is also central in our approach. This proof scheme can also be framed as a mirror descent analysis in well-chosen geometries [23, 2] and has been recently extended to non-compact settings [18] or [11, Proposition 4].
@@ 1.2 Paper outline
The rest of the paper is organized as follows. In Section 2 we define the objects of interest and recall a few useful formulas. The full statements of our main results are in Section 3, and the core of the proofs can be found in Sections 4 and 5. The convergence rate of Sinkhorn’s algorithm for isotropic Gaussian is studied in Section 6. Finally we gather in the Appendix the proof of intermediate technical results.
@@ 2.1 General conventions
$\mathrm{KL}\left(\rho|\rho^{\prime}\right)=\begin{cases}\int\frac{\mathrm{d}\rho}{\mathrm{d}\rho^{\prime}}\mathrm{d}\rho&\text{ if }\rho\ll\rho^{\prime},\\ +\infty&\text{ otherwise.}\end{cases}$ (6)
@@ 2.2 $(c,\lambda)$-transforms and the Sinkhorn algorithm
For a cost function $c$ and a regularization parameter $\lambda>0$, the $(c,\lambda)$-transforms $\phi^{c,\lambda}$,$\psi^{c,\lambda}$ of functions $\phi\in L^{1}(\mu),\psi\in L^{1}(\nu)$ are respectively defined as
@@
If $\left\|c\right\|_{\mathrm{osc}}=c_{\infty}<\infty$ (resp. $\mathrm{Lip}(c)<\infty)$, then for any $\phi\in L^{1}(\mu)$ and $\psi\in L^{1}(\nu)$,
@@
The $(c,\lambda)$-transform can be seen as an entropic equivalent of the $c$-transform in the unregularized optimal transport theory. Indeed, it can be shown by computing the first variation of the dual objective (3) that for any $\phi\in L^{1}(\mu)$ and $\psi\in L^{1}(\nu)$, we have
$F(\psi^{c,\lambda},\psi)=\max_{\phi\in L^{1}(\mu)}F(\phi,\psi)\quad\text{and}\quad F(\phi,\phi^{c,\lambda})=\max_{\psi\in L^{1}(\nu)}F(\phi,\psi).$ (12)
@@
A direct computation shows that a maximizing pair $(\phi^{*},\psi^{*})$ for the dual objective $F:L^{1}(\mu)\times L^{1}(\nu)\to\mathbb{R}$ can be obtained by solving the Schrödinger system:
@@
In particular, a solution $(\phi,\psi)$ to the Schrödinger system verifies $\phi=\psi^{c,\lambda}$ and $\psi=\psi^{\overline{c,\lambda}}$. Under our assumptions, such a maximizing pair always exists and it is unique up to translations by a constant. Hence, we may now more formally define the Sinkhorn updates initialized at an arbitrary $\psi_{0}\in L^{1}(\nu)$ by
@@
Because for every $t\geq 0$ the iterate $(\phi_{t+1/2},\psi_{t})=(\psi_{t}^{c,\lambda},\psi_{t})$ depends only on $\psi_{t}$, it will be convenient to rewrite the dual objective in terms of the function $\psi\in L^{1}(\nu)$. This leads to the notion of semi-dual objective described in the next section.
@@ 2.3 Semi-dual objective and primal-dual relations
The semi-dual objective $E:L^{1}(\nu)\to\mathbb{R}$ is defined by
$E(\psi)=\max_{\phi\in L^{1}(\mu)}F(\phi,\psi)=F(\psi^{c,\lambda},\psi).$ (16)
The definition of the $(c,\lambda)$-transforms implies that for any $\psi\in L^{1}(\nu)$
@@
Hence, for any $\psi\in L^{1}(\nu)$ the semi-dual objective $E$ can be written as
@@
Any $\psi\in L^{1}(\nu)$ induces a probability measure $\gamma[\psi]\in\mathcal{P}(\mathcal{X}\times\mathcal{Y})$ defined as
@@
The measure $\gamma[\psi]$ is, in general, not primal feasible as it does not belong to the set of couplings $\Gamma(\mu,\nu)$. In fact, we have that $\gamma[\psi]$ belongs to $\Gamma(\mu,\nu)$ if and only if $\psi$ is equal to a maximizer $\psi^{*}$ of the semi-dual objective $E$, in which case the measure $\gamma[\psi]=\gamma[\psi^{*}]$ is equal to the unique minimizer of the primal problem (1). Moreover, for any $\psi\in L^{1}(\nu)$ the following relation connects the dual sub-optimality gap to a form of a primal guarantee:
@@ 2.4 Basic properties of Sinkhorn’s iterates
We now list some basic properties satisfied by Sinkhorn’s iterates $(\phi_{t+1/2},\psi_{t})_{t\geq 0}$. They are well-known in the literature, and we will use them without explicit references to this section. For any $\psi\in L^{1}(\nu)$, define the marginal measures $\mu[\psi]\in\mathcal{P}(\mathcal{X})$ and $\nu[\psi]\in\mathcal{P}(\mathcal{Y})$ of $\gamma[\psi]\in\mathcal{P}(\mathcal{X}\times\mathcal{Y})$ by
@@
The definition of $(c,\lambda)$-transforms ensures that $\mu[\psi]=\mu$. However, $\nu[\psi]\neq\nu$ unless $\psi$ is a maximizer of the semi-dual objective $E$. An explicit calculation shows that for any $\psi\in L^{1}(\nu)$,
@@ 2.5 Entropic Kantorovich functional
We end this background section with the introduction of what we refer to as the entropic Kantorovich functional, denoted $\mathcal{K}:L^{1}(\nu)\to\mathbb{R}$ and defined by
@@
Notice that this functional corresponds to the non-linear part of the semi-dual functional $E(\psi)=\langle\psi^{c,\lambda},\mu\rangle+\langle\psi,\nu\rangle=\mathcal{K}(\rho)+\langle\psi,\nu\rangle$. The directional derivatives of $\mathcal{K}$ can be made explicit. First, for any $\psi\in L^{1}(\nu)$ and any $x\in\mathcal{X}$, we define the probability measure $\nu_{x}[\psi]$ on $\mathcal{Y}$ as
@@
The functional $\mathcal{K}$ is concave on $L^{1}(\nu)$, and it is strictly concave up to translations. Moreover, $\mathcal{K}$ admits the following first and second order directional derivatives: for any functions $\psi,v\in L^{1}(\nu)$ and any $\varepsilon\in\mathbb{R}$,
@@ 3 Main result: exponential convergence with robust contraction constants
Assume that (A) holds and that either the cost $c$ is Lipschitz continuous or the target measure $\nu$ is finitely supported. Then, there exists an integer $T$ depending on $\mu,\nu,c$ and $\psi_{0}$ such that for any $t\geq T$, guarantee (28) of Theorem 3.1 holds with a value of $\alpha$ improved to:
@@
The semi-concavity and semi-convexity assumptions made on $c$ are easily satified whenever $c$ is smooth and $\mathcal{X}$ and $\mathcal{Y}$ are compact, since in this case $c$ is both $\left\|c\right\|_{\mathcal{C}^{2}(\mathcal{X}\times\mathcal{Y})}$-semi-concave and $\left\|c\right\|_{\mathcal{C}^{2}(\mathcal{X}\times\mathcal{Y})}$-semi-convex.
@@
The statement of Corollary 3.2 is asymptotic. However, as detailed in Section B, the value of the integer $T$ present in this statement can be made explicit in some cases. For instance, whenever the target measure $\nu$ is finitely supported over $\mathcal{Y}=\{y_{1},\dots,y_{N}\}$, we have
@@
By Bernstein’s moment generating function bound for bounded random variables [29, Proposition 2.14] and the fact that $\psi^{*}-\psi_{t}$ and hence $f$ is contained in an interval of length at most $2c_{\infty}$, it follows that for any $\eta\in(0,\frac{3}{2c_{\infty}})$ we have
$\log\mathbb{E}_{y\sim\nu[\psi_{t}]}\left[\exp\left(\eta(\psi^{*}(y)-\psi_{t}(y))\right)\right]\leq\frac{\eta^{2}\mathbb{V}\mathrm{ar}_{\nu[\psi_{t}]}(\psi^{*}-\psi_{t})}{2(1-\eta\frac{2c_{\infty}}{3})}.$ (41)
@@
The above strong-concavity estimate and its proof idea relying on the Prékopa-Leindler inequality already appeared in [13, Theorem 3.2], but only in the case where $c$ was the linear ground cost $c(x,y)=-\langle x,y\rangle$. Interestingly, letting the regularization strength $\lambda$ go to 00 in the above estimate allows to also get an estimation of the strong-concavity of the non-regularized Kantorovich functional, which is the functional
@@
where $\psi^{c}(\cdot)=\inf_{y}c(\cdot,y)-\psi(y)$ is the usual $c$-transform. This allows to extend to bounded semi-concave costs the strong-concavity estimate that was already established in [14, Theorem 2.1] for the non-regularized Kantorovich functional with linear ground cost $c(x,y)=-\langle x,y\rangle$, which in turn was proven relying on the Brascamp-Lieb concentration inequality.
@@
The idea of this definition comes from [22, Exercise 2.2.11]. From Lemma A.1, the mapping $s\mapsto I_{\rho}(\psi+sv)$ is $\mathcal{C}^{2}$ on $\mathbb{R}$. Denote $\tilde{\mu}[\psi]$ the probability measure in $\mathcal{P}(\mathcal{X})$ satisfying
@@
Let $\xi>0$ and $\rho$ be a measure on $\mathbb{R}^{d}$ of the form $\mathrm{d}\rho=e^{-W}$ where $W$ is a $\xi$-strongly-convex function. Let $\alpha\in[0,1]$ and let $f,g,h:\mathbb{R}^{d}\to\mathbb{R}_{+}$ be such that for all $x,y\in\mathbb{R}^{d}$,
$h((1-\alpha)x+\alpha y)\geq e^{-\xi\alpha(1-\alpha)\left\|x-y\right\|^{2}/2}f(x)^{1-\alpha}g(y)^{\alpha}.$
@@
Let $\psi^{0},\psi^{1}:\mathbb{R}^{d}\to\mathbb{R}$, $\alpha\in(0,1)$, and assume (A). Then, denoting $\psi^{\alpha}=(1-\alpha)\psi^{0}+\alpha\psi^{1}$, it holds for any $u,v\in\mathcal{X}$ that
$(\psi^{\alpha})^{c,\lambda}((1-\alpha)u+\alpha v)\geq-\frac{\xi}{2}\alpha(1-\alpha)\left\|u-v\right\|^{2}+(1-\alpha)(\psi^{0})^{c,\lambda}(u)+\alpha(\psi^{1})^{c,\lambda}(v).$ (89)
@@
$c((1-\alpha)u+\alpha v,y)-\frac{\xi}{2}\left\|(1-\alpha)u+\alpha v\right\|^{2}\geq(1-\alpha)(c(u,y)-\frac{\xi}{2}\left\|u\right\|^{2})+\alpha(c(v,y)-\frac{\xi}{2}\left\|v\right\|^{2}).$
@@
$c((1-\alpha)u+\alpha v,y)\geq(1-\alpha)c(u,y)+\alpha c(v,y)-\frac{\xi}{2}\alpha(1-\alpha)\left\|u-v\right\|^{2}.$

### v2

@@ Abstract
We study the convergence rate of Sinkhorn’s algorithm for solving entropy-regularized optimal transport problems when at least one of the probability measures, $\mu$, admits a density over $\mathbb{R}^{d}$. For a semi-concave cost function bounded by $c_{\infty}$ and a regularization parameter $\lambda>0$, we obtain exponential convergence guarantees on the dual sub-optimality gap with contraction rates that are polynomial in $\lambda/c_{\infty}$. This represents an exponential improvement over the known contraction rate $1-\Theta(\exp(-c_{\infty}/\lambda))$ achievable via Hilbert’s projective metric. Specifically, we prove a contraction rate value of $1-\Theta(\lambda^{2}/c_{\infty}^{2})$ when $\mu$ has a bounded log-density. In some cases, such as when $\mu$ is log-concave and the cost function is $c(x,y)=-\langle x,y\rangle$, this rate improves to $1-\Theta(\lambda/c_{\infty})$. The latter rate matches the one that we derive for the transport between isotropic Gaussian measures, indicating tightness in the dependency in $\lambda/c_{\infty}$. Our results are fully non-asymptotic and explicit in all the parameters of the problem.
Keywords — Sinkhorn’s algorithm, entropic optimal transport, Schrödinger bridge problem, linear convergence.
2020 Mathematics Subject Classification — 90C08, 90-08, 65K10, 49M29.
@@ 1 Introduction
$\sup_{\phi\in L^{\infty}(\mu),\psi\in L^{\infty}(\nu)}\int\phi\mathrm{d}\mu+\int\psi\mathrm{d}\nu+\lambda\left(1-\int\int\exp\left(\frac{\phi\oplus\psi-c}{\lambda}\right)\mathrm{d}\mu\mathrm{d}\nu\right),$ (3)
where $(\phi\oplus\psi)(x,y)=\phi(x)+\psi(y)$. The classical approach to solve (3) is via Sinkhorn’s algorithm. Denoting $F(\phi,\psi)$ the dual functional being maximized in (3), Sinkhorn’s algorithm starts at an arbitrary point $\psi_{0}\in L^{\infty}(\nu)$ and repeatedly performs exact maximization of the objectives $\phi\mapsto F(\phi,\psi)$ and $\psi\mapsto F(\phi,\psi)$. This leads to the sequence of iterates
@@ 1.1 Comparison with related work
Iterations (4), which we refer to as Sinkhorn’s algorithm [37], were rediscovered several times and perhaps first appeared in [39]. Consequently, these iterations and close variants thereof are known under various names in various communities, such as the Sinkhorn-Knopp algorithm, Iterative Proportional Fitting Procedure (IPFP), RAS algorithm, matrix scaling algorithm, Bregman algernative projection or Fortet’s iterations; see the surveys [26, 33] for some historical remarks. Most quantitative convergence analyses can be classified into two categories: (i) exponential convergence rates with non-robust constants and (ii) polynomial convergence rates with robust constants.
Convergence guarantees of the first kind can be obtained by noticing that Sinkhorn’s iteration is a contraction for Hilbert’s projective metric [22, 5] where the contraction rate can be shown to be at most $1-\Theta(e^{-c_{\infty}/\lambda})$; see [10, 19] for extensions to continuous settings and [21] for non-compact settings, as well as [33, Remark 4.12] for a bibliographical discussion. Note that this type of contraction rate can also be obtained without using Hilbert’s projective metric, for example by using a classical convex optimization approach as in [32, 7], allowing to extend the analysis to multi-marginal settings; or by relying on coupling methods and stochastic control following [25, 13], allowing extensions to non-bounded configurations. As put forth in [33, Remark 4.15], the dependence on $c_{\infty}/\lambda$ of the $1-\Theta(e^{-c_{\infty}/\lambda})$ contraction rates numerically seems tight in the most difficult cases, e.g. when the cost $c$ is random. In more favorable cases however, such as when there exists an optimal transport map solving problem (1) with $\lambda=0$, many have observed experimentally that the dependency in $c_{\infty}/\lambda$ is generally much better in practice.
Convergence guarantees of the second kind can be traced back to [28], and have more recently regained interest [1, 9, 20]. In their tightest form, due to [20], one has the guarantee $\delta_{t}\leq 2c_{\infty}^{2}/(\lambda t)$ for $t\geq 1$ (see [12, Proposition 10]). This result relies on an explicit expression (see Eq. (25)) for the one-step dual improvement $\delta_{t}-\delta_{t+1}$, which is also central in our approach. This proof scheme can also be framed as a mirror descent analysis in well-chosen geometries [30, 2] and has been recently extended to non-compact settings [24] or [15, Proposition 4].
@@ 1.2 Paper outline
The rest of the paper is organized as follows. In Section 2 we define the objects of interest and recall a few useful formulas. The full statements of our main results are in Section 3, which also contains an outline of the proofs in Section 3.1. The core of the proofs can be found in Sections 4 and 5. The convergence rate of Sinkhorn’s algorithm for isotropic Gaussian is studied in Section 6. Section 7 gathers numerical experiments indicating that the core feature of our theoretical findings also holds in practical settings where both marginals are discrete probability distributions. Finally we gather in the Appendix the proof of intermediate technical results.
@@ 2.1 General conventions
$\mathrm{KL}\left(\rho|\rho^{\prime}\right)=\begin{cases}\int\log\frac{\mathrm{d}\rho}{\mathrm{d}\rho^{\prime}}\mathrm{d}\rho&\text{ if }\rho\ll\rho^{\prime},\\ +\infty&\text{ otherwise.}\end{cases}$ (6)
@@ 2.2 $(c,\lambda)$-transforms and the Sinkhorn algorithm
For a cost function $c$ and a regularization parameter $\lambda>0$, the $(c,\lambda)$-transforms $\phi^{c,\lambda}$,$\psi^{c,\lambda}$ of functions $\phi\in L^{\infty}(\mu),\psi\in L^{\infty}(\nu)$ are respectively defined as
@@
If $\left\|c\right\|_{\mathrm{osc}}=c_{\infty}<\infty$ (resp. $\mathrm{Lip}(c)<\infty)$, then for any $\phi\in L^{\infty}(\mu)$ and $\psi\in L^{\infty}(\nu)$,
@@
The $(c,\lambda)$-transform can be seen as an entropic equivalent of the $c$-transform in the unregularized optimal transport theory. Indeed, it can be shown by computing the first variation of the dual objective (3) that for any $\phi\in L^{\infty}(\mu)$ and $\psi\in L^{\infty}(\nu)$, we have
$F(\psi^{c,\lambda},\psi)=\max_{\phi\in L^{\infty}(\mu)}F(\phi,\psi)\quad\text{and}\quad F(\phi,\phi^{c,\lambda})=\max_{\psi\in L^{\infty}(\nu)}F(\phi,\psi).$ (12)
@@
A direct computation shows that a maximizing pair $(\phi^{*},\psi^{*})$ for the dual objective $F:L^{\infty}(\mu)\times L^{\infty}(\nu)\to\mathbb{R}$ can be obtained by solving the Schrödinger system:
@@
In particular, a solution $(\phi,\psi)$ to the Schrödinger system verifies $\phi=\psi^{c,\lambda}$ and $\psi=\psi^{\overline{c,\lambda}}$. Under our assumptions, such a maximizing pair always exists and it is unique up to translations by a constant. Hence, we may now more formally define the Sinkhorn updates initialized at an arbitrary $\psi_{0}\in L^{\infty}(\nu)$ by
@@
Because for every $t\geq 0$ the iterate $(\phi_{t+1/2},\psi_{t})=(\psi_{t}^{c,\lambda},\psi_{t})$ depends only on $\psi_{t}$, it will be convenient to rewrite the dual objective in terms of the function $\psi\in L^{\infty}(\nu)$. This leads to the notion of semi-dual objective described in the next section.
@@ 2.3 Semi-dual objective and primal-dual relations
The semi-dual objective $E:L^{\infty}(\nu)\to\mathbb{R}$ is defined by
$E(\psi)=\max_{\phi\in L^{\infty}(\mu)}F(\phi,\psi)=F(\psi^{c,\lambda},\psi).$ (16)
The definition of the $(c,\lambda)$-transforms implies that for any $\psi\in L^{\infty}(\nu)$
@@
Hence, for any $\psi\in L^{\infty}(\nu)$ the semi-dual objective $E$ can be written as
@@
Any $\psi\in L^{\infty}(\nu)$ induces a probability measure $\gamma[\psi]\in\mathcal{P}(\mathcal{X}\times\mathcal{Y})$ defined as
@@
The measure $\gamma[\psi]$ is, in general, not primal feasible as it does not belong to the set of couplings $\Gamma(\mu,\nu)$. In fact, we have that $\gamma[\psi]$ belongs to $\Gamma(\mu,\nu)$ if and only if $\psi$ is equal to a maximizer $\psi^{*}$ of the semi-dual objective $E$, in which case the measure $\gamma[\psi]=\gamma[\psi^{*}]$ is equal to the unique minimizer of the primal problem (1). Moreover, for any $\psi\in L^{\infty}(\nu)$ the following relation connects the dual sub-optimality gap to a form of a primal guarantee:
@@ 2.4 Basic properties of Sinkhorn’s iterates
We now list some basic properties satisfied by Sinkhorn’s iterates $(\phi_{t+1/2},\psi_{t})_{t\geq 0}$. They are well-known in the literature, and we will use them without explicit references to this section. For any $\psi\in L^{\infty}(\nu)$, define the marginal measures $\mu[\psi]\in\mathcal{P}(\mathcal{X})$ and $\nu[\psi]\in\mathcal{P}(\mathcal{Y})$ of $\gamma[\psi]\in\mathcal{P}(\mathcal{X}\times\mathcal{Y})$ by
@@
The definition of $(c,\lambda)$-transforms ensures that $\mu[\psi]=\mu$. However, $\nu[\psi]\neq\nu$ unless $\psi$ is a maximizer of the semi-dual objective $E$. An explicit calculation shows that for any $\psi\in L^{\infty}(\nu)$,
@@ 2.5 Entropic Kantorovich functional
We end this background section with the introduction of what we refer to as the entropic Kantorovich functional, denoted $\mathcal{K}:L^{\infty}(\nu)\to\mathbb{R}$ and defined by
@@
Notice that this functional corresponds to the non-linear part of the semi-dual functional $E(\psi)=\langle\psi^{c,\lambda},\mu\rangle+\langle\psi,\nu\rangle=\mathcal{K}(\psi)+\langle\psi,\nu\rangle$. The directional derivatives of $\mathcal{K}$ can be made explicit. First, for any $\psi\in L^{\infty}(\nu)$ and any $x\in\mathcal{X}$, we define the probability measure $\nu_{x}[\psi]$ on $\mathcal{Y}$ as
@@
The functional $\mathcal{K}$ is concave on $L^{\infty}(\nu)$, and it is strictly concave up to translations. Moreover, $\mathcal{K}$ admits the following first and second order directional derivatives: for any functions $\psi,v\in L^{\infty}(\nu)$ and any $\varepsilon\in\mathbb{R}$,
@@ 3 Main result: exponential convergence with robust contraction constants
Assume that (A) holds and that either the cost $c$ is Lipschitz continuous and $\mathcal{Y}$ is bounded or the target measure $\nu$ is finitely supported. Then, there exists an integer $T$ depending on $\mu,\nu,c$ and $\psi_{0}$ such that for any $t\geq T$, guarantee (28) of Theorem 3.1 holds with a value of $\alpha$ improved to:
@@
The semi-concavity and semi-convexity assumptions made on $c$ are easily satisfied whenever $c$ is smooth and $\mathcal{X}$ and $\mathcal{Y}$ are compact, since in this case $c$ is both $\left\|c\right\|_{\mathcal{C}^{2}(\mathcal{X}\times\mathcal{Y})}$-semi-concave and $\left\|c\right\|_{\mathcal{C}^{2}(\mathcal{X}\times\mathcal{Y})}$-semi-convex.
Remark 3.3.
Assumptions ($A_{1}$), ($A_{2}$) or ($A_{3}$) of Theorem 3.1 all require the measure $\mu$ to be absolutely continuous. Because of this assumption, the $(c,\lambda)$-transform
$\phi^{c,\lambda}(y)=-\lambda\log\int e^{\frac{\phi(x)-c(x,y)}{\lambda}}\mathrm{d}\mu(x)$
appearing in the definition of Sinkhorn’s iterates (15) cannot be implemented exactly on a computer in general. In this sense, the convergence result of Theorem 3.1 concerns an infinite dimensional or continuous in space version of Sinkhorn’s algorithm. We notice however in Section C of the appendix that the results of Theorem 3.1 can be extended to the practical setting where the above $(c,\lambda)$-transform is only approximated. Let us consider for instance that we access a discrete approximation $\hat{\mu}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}$ of $\mu$ that is such that for any potential $\phi\in L^{\infty}(\mu)$ having the same regularity as the cost function $c$,
$\left\|\phi^{\widehat{c,\lambda}}-\phi^{c,\lambda}\right\|_{\infty}\leq\varepsilon$
for some fixed $\varepsilon>0$, where $\phi^{\widehat{c,\lambda}}(y)=-\lambda\log\frac{1}{N}\sum_{i=1}^{N}e^{\frac{\phi(x_{i})-c(x_{i},y)}{\lambda}}$. Then assuming (A) and either ($A_{1}$) or ($A_{2}$) for $\mu$ (no further assumption is made on $\hat{\mu}$), Proposition C.1 ensures that that the approximated Sinkhorn iterates $\hat{\psi}_{t+1}=(\hat{\psi}_{t}^{c,\lambda})^{\widehat{c,\lambda}}$ satisfy
$E(\psi^{*})-E(\hat{\psi}_{t})\leq(1-\alpha^{-1})^{t}(E(\psi^{*})-E(\hat{\psi}_{0}))+2\alpha\varepsilon.$
This corresponds to an exponential contraction of the dual population objective $E$ (that is associated to $\mu$), up to an error term which is proportional to the error made in the approximation of the $(c,\lambda)$-transforms.
@@
The statement of Corollary 3.2 is asymptotic. However, as detailed in Section B of the Appendix, the value of the integer $T$ present in this statement can be made explicit in some cases. For instance, whenever the target measure $\nu$ is finitely supported over $\mathcal{Y}=\{y_{1},\dots,y_{N}\}$, we have
@@
By Bernstein’s moment generating function bound for bounded random variables [38, Proposition 2.14] and the fact that $\psi^{*}-\psi_{t}$ and hence $f$ is contained in an interval of length at most $2c_{\infty}$, it follows that for any $\eta\in(0,\frac{3}{2c_{\infty}})$ we have
$\log\mathbb{E}_{y\sim\nu[\psi_{t}]}\left[\exp\left(\eta f\right)\right]\leq\frac{\eta^{2}\mathbb{V}\mathrm{ar}_{\nu[\psi_{t}]}(\psi^{*}-\psi_{t})}{2(1-\eta\frac{2c_{\infty}}{3})}.$ (41)
@@
The above strong-concavity estimate and its proof idea relying on the Prékopa-Leindler inequality already appeared in [17, Theorem 3.2], but only in the case where $c$ was the linear ground cost $c(x,y)=-\langle x,y\rangle$. Interestingly, letting the regularization strength $\lambda$ go to 00 in the above estimate allows to also get an estimation of the strong-concavity of the non-regularized Kantorovich functional, which is the functional
@@
where $\psi^{c}(\cdot)=\inf_{y}c(\cdot,y)-\psi(y)$ is the usual $c$-transform. This allows to extend to bounded semi-concave costs the strong-concavity estimate that was already established in [18, Theorem 2.1] for the non-regularized Kantorovich functional with linear ground cost $c(x,y)=-\langle x,y\rangle$, which in turn was proven relying on the Brascamp-Lieb concentration inequality.
@@
The idea of this definition comes from [29, Exercise 2.2.11]. From Lemma A.1, the mapping $s\mapsto I_{\rho}(\psi+sv)$ is $\mathcal{C}^{2}$ on $\mathbb{R}$. Denote $\tilde{\mu}[\psi]$ the probability measure in $\mathcal{P}(\mathcal{X})$ satisfying
@@
Let $\xi>0$ and $\rho$ be a measure on $\mathbb{R}^{d}$ of the form $\mathrm{d}\rho=e^{-W}$ where $W$ is a $\xi$-strongly-convex function. Let $s\in[0,1]$ and let $f,g,h:\mathbb{R}^{d}\to\mathbb{R}_{+}$ be such that for all $x,y\in\mathbb{R}^{d}$,
$h((1-s)x+sy)\geq e^{-\xi s(1-s)\left\|x-y\right\|^{2}/2}f(x)^{1-s}g(y)^{s}.$
@@
Let $\psi^{0},\psi^{1}:\mathbb{R}^{d}\to\mathbb{R}$, $s\in(0,1)$, and assume (A). Then, denoting $\psi^{s}=(1-s)\psi^{0}+s\psi^{1}$, it holds for any $u,v\in\mathcal{X}$ that
$(\psi^{s})^{c,\lambda}((1-s)u+sv)\geq-\frac{\xi}{2}s(1-s)\left\|u-v\right\|^{2}+(1-s)(\psi^{0})^{c,\lambda}(u)+s(\psi^{1})^{c,\lambda}(v).$ (89)
@@
$c((1-s)u+sv,y)-\frac{\xi}{2}\left\|(1-s)u+sv\right\|^{2}\geq(1-s)(c(u,y)-\frac{\xi}{2}\left\|u\right\|^{2})+s(c(v,y)-\frac{\xi}{2}\left\|v\right\|^{2}).$
@@
$c((1-s)u+sv,y)\geq(1-s)c(u,y)+sc(v,y)-\frac{\xi}{2}s(1-s)\left\|u-v\right\|^{2}.$